The CCU manager is to report the inpatient service days to the director of nursing each day. On November 15, the CCU had 22 patients at 12:01 a.m. The CCU received one new direct admission, one transfer-in from 3 West, and one transfer-in from 2 East. The CCU transferred three patients out to 3 West, discharged one to a hospice center across town, and had one patient die on November 15.

What are the inpatient service days for the CCU on November 15? Calculate the total inpatient service days for Dover Regional Medical Center for the CCU on November 15.
Answer #1

Answer:

Inpatient service days is a unit of measure equivalent to the services received by one inpatient during one 24- hour period.

census = 22   

admission = 3

22 + 3 = 25

Discharge = 4

25 - 4 = 21

21 - 1 died = 20 total Inpatient services days.